OpenStudy (dariyana):

What is the m∠ADC? 14 24 44 134

hero (hero):

What the...

hero (hero):

First you were orange, now you're yellow

OpenStudy (dariyana):

lol iam cool like tht

OpenStudy (dariyana):

jk @Hero

OpenStudy (anonymous):

m∠DCA and m∠DCB are supplementary angles. Therefore: m∠DCA + m∠DCB = 180 m∠DCB = 46, solve for m∠DCA. Once you've done that you'll have the values of two of the three angles in the triangle. At that point you can solve for m∠ADC. m∠DCA + m∠DAC + m∠ADC = 180

OpenStudy (anonymous):

BTW if it gets confusing with all of the angles with similar names, just rewrite them like so: m∠DCB = W = 46 m∠DCA = X m∠DAC = Y = 22 m∠ADC = Z 180 = W + X = 46 + X 180 = X + Y + Z = X + 22 + Z

OpenStudy (dariyana):

i dont get it so what would answer be and how @Wired

OpenStudy (anonymous):

W = 46 180 = W + X Solve for X.

OpenStudy (dariyana):

so 134

OpenStudy (anonymous):

Yep. X = 134 Y = 22 180 = X + Y + Z

OpenStudy (dariyana):

thank you